OpenStudy (anonymous):

some one please help me solve this problem find the inverse of y=7x2-3

OpenStudy (anonymous):

x=7y^2-3 x+3=7y^2 x+3/7=y^2 y=square root of x+3/7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np